OpenStudy (anonymous):

If f(x) = tan^(2)x + sinx, then f'(pi/4) =

OpenStudy (callisto):

Do you know how to find f'(x)?

OpenStudy (anonymous):

would it be \[f'(x)=2tanxsec^2x+cosx\] ?

OpenStudy (anonymous):

would it? :o

OpenStudy (ash2326):

that's right

OpenStudy (anonymous):

Okay i just have issues with the rest when evaluating f'(pi/4)

OpenStudy (callisto):

What are tan(pi/4) and cos (pi/4)?

OpenStudy (anonymous):

tan(pi/4)=1 and cos(pi/4)=\[\frac{ 1 }{ \sqrt{2} }\]

OpenStudy (callisto):

Nice :) What is sec(pi/4)? Note that sec(theta) = 1/[cos(theta)]

OpenStudy (anonymous):

so that would equal to 2?

OpenStudy (callisto):

sec(pi/4) = 1/cos(pi/4) = \(\large \frac{1}{\frac{1}{\sqrt{2}}}\) = \(\sqrt{2}\) So, sec^2 (pi/4) = 2. Does that make sense to you?

OpenStudy (anonymous):

yes i understand that, so now how would the final answer look put together

OpenStudy (callisto):

f'(x) = 2tanx sec^2(x) + cosx f'(pi/4) = 2 tan(pi/4) sec^2(pi/4) + cos(pi/4) Just now, you have evaluated tan(pi/4), sec^2(pi/4) and cos(pi/4), just substitute the values you got into f'(pi/4), so you get \[f'(\frac{\pi}{4}) = 2 tan(\frac{\pi}{4}) sec^2(\frac{\pi}{4}) + cos(\frac{\pi}{4})\] \[f'(\frac{\pi}{4}) = 2(1)(2)+\frac{1}{\sqrt{2}}=...???\]

OpenStudy (anonymous):

would it be \[\frac{ 4\sqrt{2}+1 }{\sqrt{2} }\]

mathslover (mathslover):

Yeah! @ifunfrank it is \(\cfrac { 4\sqrt{2} + 1 }{\sqrt{2} }\) Good work!

mathslover (mathslover):

You can also rationalize it : \(\cfrac{4\sqrt{2}+1}{\sqrt{2}} \times \cfrac{\sqrt{2}}{\sqrt{2}}\) \(\implies \cfrac{8 + \sqrt{2}}{2}\)